The length of the base of an isosceles triangle is 10 meters less than twice the length of one of its legs. If the perimeter of the triangle is 50 meters, find the length of the base of the triangle.
Answer by lwsshak3(11628)

The length of the base of an isosceles triangle is 10 meters less than twice the length of one of its legs. If the perimeter of the triangle is 50 meters, find the length of the base of the triangle.
**
let x=length of one of the legs
2x-10=length of the base
Perimeter =50=2x+2x-10
4x=60
x=15
2x-10=30-10=20
ans:
Base=20 meters